# AntibodyDesignBFN: High-Fidelity Fixed-Backbone Antibody Design via Discrete Bayesian Flow Networks

### v1

@@ Abstract
The computational design of antibodies with high specificity and affinity is a cornerstone of modern therapeutic development. While deep generative models, particularly Denoising Diffusion Probabilistic Models (DDPMs), have demonstrated the ability to generate realistic antibody structures, they often suffer from high computational costs and the difficulty of modeling discrete variables like amino acid sequences. In this work, we present AntibodyDesignBFN, a novel framework for fixed-backbone antibody design based on Discrete Bayesian Flow Networks (BFN). Unlike standard diffusion models that rely on Gaussian noise removal or complex discrete corruption processes, BFNs operate directly on the parameters of the data distribution, enabling a continuous-time, fully differentiable generative process on the probability simplex. While recent pioneering works like IgCraft and AbBFN have introduced BFNs to the domain of antibody sequence generation and inpainting, our work focuses specifically on the inverse folding task—designing sequences that fold into a fixed 3D backbone. By integrating a lightweight Geometric Transformer utilizing Invariant Point Attention (IPA) and a resource-efficient training strategy with gradient accumulation, our model achieves superior performance. Evaluations on a rigorous 2025 temporal test set reveal that AntibodyDesignBFN achieves a robust 49.9% Overall Amino Acid Recovery (AAR), demonstrating that BFNs, when conditioned on 3D geometric constraints, offer a robust mathematical framework for high-fidelity antibody design. Code and model checkpoints are available at https://github.com/YueHuLab/AntibodyDesignBFN and https://huggingface.co/YueHuLab/AntibodyDesignBFN, respectively.

### v2

@@ Abstract
The computational design of antibodies with high specificity and affinity is a cornerstone of modern therapeutic development. While deep generative models, particularly Denoising Diffusion Probabilistic Models (DDPMs), have demonstrated the ability to generate realistic antibody structures, they often suffer from high computational costs and the difficulty of modeling discrete variables like amino acid sequences. In this work, we present AntibodyDesignBFN, a novel framework for fixed-backbone antibody design based on Discrete Bayesian Flow Networks (BFN). Unlike standard diffusion models that rely on Gaussian noise removal or complex discrete corruption processes, BFNs operate directly on the parameters of the data distribution, enabling a continuous-time, fully differentiable generative process on the probability simplex. While recent pioneering works like IgCraft and AbBFN have introduced BFNs to the domain of antibody sequence generation and inpainting, our work focuses specifically on the inverse folding task—designing sequences that fold into a fixed 3D backbone. By integrating a lightweight Geometric Transformer utilizing Invariant Point Attention (IPA) and a resource-efficient training strategy with gradient accumulation, our model achieves superior performance. Evaluations on a rigorous 2025 temporal test set reveal that AntibodyDesignBFN achieves superior Amino Acid Recovery (AAR) compared to strong baselines, demonstrating that BFNs, when conditioned on 3D geometric constraints, offer a robust mathematical framework for high-fidelity antibody design. Code and model checkpoints are available at https://github.com/YueHuLab/AntibodyDesignBFN and https://huggingface.co/YueHuLab/AntibodyDesignBFN, respectively.